# Pointwise convergence to initial data of heat and Hermite-heat equations in Modulation Spaces

### v4

@@ Abstract.
We characterize weighted modulation spaces (data space) for which the heat semigroup $e^{-tL}f$ converges pointwise to the initial data $f$ as time $t$ tends to zero. Here $L$ stands for the standard Laplacian $\Delta$ or Hermite operator $H=-\Delta+|x|^{2}$ on the Euclidean space. This is the first result on pointwise convergence with data in a weighted modulation spaces (which do not coincide with weighted Lebesgue spaces). We also prove that the Hardy-Littlewood maximal operator operates on certain modulation spaces. This may be of independent interest.
@@ 1.1. Background and motivation
The aim of this paper to study under what condition on initial data $f,$ the solution to (1.1) converges to $f$ pointwise? Specifically, we wish to characterize the class of data for which the following limit
@@
where $h^{H}_{t}(x,y)$ denotes the corresponding heat kernel associated with $H$. See Section 6 for details. The pointwise convergence problem is closely connected with the boundendess of Hardy–Littlewood maximal operator $\mathcal{M}f$, which is defined by
@@
$D^{h}_{p}(\mathbb{R}^{n}):=\left\{v:\mathbb{R}^{n}\to(0,\infty):\exists\,t_{0}>0\ni\ h_{t_{0}}\in L^{p^{\prime}}_{v^{-1}}(\mathbb{R}^{n})\right\}$ (1.7)
@@
    There exists $0<R<\infty$ such that $W_{R}*f(x)<\infty$ a.e. $x$ and the limit $\lim_{t\rightarrow 0}h_{t}*f(x)$ exists a.e. $x$ for all $f\in L^{p}_{v}$.
@@
We conclude this section by highlighting the following remark. In all the aforementioned results, the the characterized class of weights and initial data are based on $L^{p}$ spaces. On the other hand, since the heat semigroup has been extensively studied on several other function spaces, it is natural to study the a.e. convergence problem there. The aim of this paper is to study this problem in the realm of modulation spaces.
@@ 1.2. The weight-class in the framework of modulation spaces
    The concept of a weight function arises frequently in harmonic analysis, which serves to quantify growth, decay, or smoothness. It turns out that the correct class of weight function to be considered is the “sub multiplicative weight” to study the weighted modulation spaces. This contrasts with the weights appearing in the classical setting, e.g. the Muckenhoupt classes. The main drawback with the Muckenhoupt classes is their lack of invariance under time–frequency shifts in weighted modulation spaces. We refer to excellent survey article by Gröchenig [27] for a most relavent classes of weights appearing in the time-frequency analysis.
@@
$D^{h}_{p,q}(\mathbb{R}^{2n})=\left\{v:\mathbb{R}^{2n}\to\mathbb{(}0,\infty):\exists\,t_{0}>0\ni h_{t_{0}}\in M^{p^{\prime},q^{\prime}}_{v^{-1}}(\mathbb{R}^{n})\right\}.$
@@
    The class of Muckenhoupt weights $A_{p}$ is a proper subset of the weight class $D_{p}^{h}$ appeared in Theorem $1.$
@@
    Our formulation of weight class $D^{h}_{p,q}$ (defined using modulation spaces) is inspired from the weight class $D^{h}_{p}$ (defined using Lebesgue spaces, see Theorem 1.1). This provides the natural way to introduce weighted structures in modulation spaces, and turned out very effective our analysis (see Remark 1.2 (2)).
@@ 2. Main results
    It is known (see e.g. [9, example 2.1]) that, for $0<\alpha<n,$ we have
@@
We shall now briefly comment on the proof of Theorem 2.1. These problems were motivated by the seminal work of Torrea et al. in [25], where they addressed similar issues in the setting of weighted Lebesgue spaces on $\mathbb{R}^{n}$. We observe that pointwise convergence results and the boundedness of the maximal operator are closely related. However, to the best of author’s knowledge, it is important to note that the boundedness of the maximal operator on modulation spaces remains an open problem (Section 2.1), requiring the development of novel techniques, which may be of independent interest. Specifically, to prove Theorem 2.1, we employ various properties of modulation spaces to reduce the problem to a stage where we can apply known results [1, 25, 28] on weighted Lebesgue spaces. We refer to Remark 5.1 for detail proof strategy.
@@ 2.1. Maximal operator
Theorem 2.4 remains unknown for arbitrary functions (not necessarily non-negative) within the modulation spaces. This discrepancy arises from the fact that the definition of the maximal operator involves $|f|$, while certain modulation spaces do not necessarily remain closed under the modulus operation, see e.g. [5, Corollary 1.3] and [11, Corollary 3.2].
@@
It is worth noting that, by exploiting certain algebraic properties inherent to modulation spaces, Theorem 2.4 remains valid for arbitrary functions (not necessarily non-negative) belonging to specific modulation spaces, when $f\in M^{p,q}$ with $1\leq p<\infty$ and $1\leq q\leq 2$. Consequently, the pointwise convergence results stated in Theorem 2.1 also hold for such modulation spaces.
@@ 4. Hardy-Littlewood maximal operator on modulation spaces
In the definition of modulation spaces, since the choice of $\phi\in\mathcal{S}\setminus\{0\}$ is arbitrary, we may select $\phi$ such that $\phi(y)\leq 0$ for a.e. $y\in\mathbb{R}^{n}$, and ensure that $f*|\phi|\in L^{p}$ by utilizing the following chain of embeddings
@@ 5. Heat equation with Laplacian
    By Lemma 5.2, the weight class on $\mathbb{R}^{2n}$, as considered in Theorem 2.1, can be suitably related to the previously studied studied weight class on $\mathbb{R}^{n}$ (see (1.7)). This correspondence ensures that the weight class on $\mathbb{R}^{2n}$ satisfies the necessary structural conditions required to apply existing results from the theory of weighted Lebesgue spaces.
@@
Hence, it follows that $g_{t_{0}}(x)<\infty$ a.e. for $x$, since we are given that $\left\|h_{\frac{t_{0}}{8}}\right\|_{M_{v^{-1}}^{p^{\prime},q^{\prime}}}<\infty.$ Now, we will choose a weight $u\in L_{\text{loc}}^{1}$ such that $g_{t_{0}}\in M_{u}^{p,q}$, which is equivalent to showing $V_{\phi}g_{t_{0}}\in L_{u}^{p,q}$ for some $\phi\in\mathcal{S}.$ By choosing $u\in L_{\text{loc}}^{1}$ as
@@
Conversely, let $\left\|h_{t_{0}}*f\right\|_{M_{u}^{p,q}}\lesssim\left\|f\right\|_{M_{v}^{p,q}}$ for all $f\in M_{v}^{p,q}$. Thus, $|h_{t_{0}}*f(x)|<\infty$ almost everywhere for all $f\in M_{v}^{p,q}$. Fix $x_{0}$ such that $h_{t_{0}}*f(x_{0})<\infty$. Then we will show that $h_{\frac{t_{0}}{4}}*f(x)<\infty$ for all $x$. Let assume $x\neq x_{0}$. Note that
@@ 6. Heat equation with Hermite operator
Consider the heat equation (1.1) associated with the harmonic oscillator $H$, whose heat semigroup is given by (1.4). The spectral decomposition of the Hermite operator is expressed in (1.3), where the eigenfunctions are the normalized Hermite functions $\Phi_{\alpha}(x)=\prod_{j=1}^{n}\tilde{h}_{\alpha_{j}}(x_{j})$ and each $\tilde{h}_{k}$ denotes the one–dimensional normalized Hermite function, given by $\tilde{h}_{k}(x)=\left(\sqrt{\pi},2^{k}k!\right)^{-1/2}(-1)^{k}e^{\frac{x^{2}}{2}}\frac{d^{k}}{dx^{k}}\big(e^{-x^{2}}\big).$ These functions form an orthonormal basis for $L^{2}$ and satisfy the eigenvalue relation $H\Phi_{\alpha}=(2|\alpha|+n)\Phi_{\alpha}$, where $|\alpha|=\alpha_{1}+\dots+\alpha_{n}$. A fundamental identity involving Hermite functions is Mehler’s formula (see [42, Lemma 1.1.1]), which for $|w|<1$ takes the form
@@
Conversely, let us assume $\lim_{t\rightarrow 0}e^{-tH}f(x)=f(x)$ almost everywhere for all non-negative $f\in M_{v}^{p,q}$. Then, by (6.4) there exists $t_{0}>0$ such that $e^{-tH}f(x)\lesssim h_{t_{0}}*f(x)<\infty$ almost everywhere for all non-negative $f\in M_{v}^{p,q}$. Following the same approach as in the converse part of Proposition 5.4, we can similarly deduce that $h_{\frac{t_{0}}{4}}\in M_{v^{-1}}^{p^{\prime},q^{\prime}}$, which implies $v\in D^{h}_{p,q}$. ∎
@@ Acknowledgement
The second author gratefully acknowledges the support provided by IISER Pune, Government of India.

### v5

@@ Abstract.
We characterize weighted modulation spaces (data space) for which the heat semigroup $e^{-tL}f$ converges pointwise to the initial data $f$ as time $t$ tends to zero. Here $L$ stands for the standard Laplacian $\Delta$ or Hermite operator $H=-\Delta+|x|^{2}$ on the Euclidean space. This is the first result on pointwise convergence with data in a weighted modulation spaces (which do not coincide with weighted Lebesgue spaces). We also prove that the Hardy-Littlewood maximal operator operates on certain modulation spaces. This may be of independent interest. We have highlighted several open questions that arise naturally from our findings.
@@ 1.1. Background and motivation
The aim of this paper is to study under what condition on initial data $f,$ the solution $u(x,t)$ to heat equation (1.1)-which is the heat semigroup as in (1.2) and (1.5) corresponding to Laplacian and Hermite operator respectively, converges to $f$ pointwise? Specifically, we wish to characterize the class of data for which the following limit
@@
where $h^{H}_{t}(x,y)$ denotes the corresponding heat kernel associated with $H$. See Section 6 for details. The pointwise convergence problem is closely connected with the boundedness of Hardy–Littlewood maximal operator $\mathcal{M}f$, which is defined by
@@
$D^{h}_{p}(\mathbb{R}^{n}):=\left\{v:\mathbb{R}^{n}\to(0,\infty)\mid\exists\,t_{0}>0\ \text{ such that }\ h_{t_{0}}\in L^{p^{\prime}}_{v^{-1}}(\mathbb{R}^{n})\right\}$ (1.8)
@@
    There exists $0<R<\infty$ such that $h_{R}*f(x)<\infty$ a.e. $x$ and the limit $\lim_{t\rightarrow 0}h_{t}*f(x)$ exists a.e. $x$ for all $f\in L^{p}_{v}$.
@@
We conclude this section by highlighting the following remark. In all the aforementioned results, the characterized class of weights and initial data are based on $L^{p}$ spaces. On the other hand, since the heat semigroup has been extensively studied on several other function spaces, it is natural to study the a.e. convergence problem there. The aim of this paper is to study this problem in the realm of modulation spaces.
@@ 1.2. The weight-class in the framework of modulation spaces
    The concept of a weight function arises frequently in harmonic analysis, which serves to quantify growth, decay, or smoothness. It turns out that the correct class of weight function to be considered is the “sub multiplicative weight” to study the weighted modulation spaces. This contrasts with the weights appearing in the classical setting, e.g. the Muckenhoupt classes. The main drawback with the Muckenhoupt classes is their lack of invariance under time–frequency shifts in weighted modulation spaces. We refer to excellent survey article by Gröchenig [27] for a most relevant classes of weights appearing in the time-frequency analysis.
@@
$D^{h}_{p,q}(\mathbb{R}^{2n})=\left\{v:\mathbb{R}^{2n}\to\mathbb{(}0,\infty)\mid\exists\,t_{0}>0\ \text{such that }h_{t_{0}}\in M^{p^{\prime},q^{\prime}}_{v^{-1}}(\mathbb{R}^{n})\right\}.$
@@
    The class of Muckenhoupt weights $A_{p}$ is a proper subset of the weight class $D_{p}^{h}$ appeared in Theorem 1.1.
@@
    Our formulation of weight class $D^{h}_{p,q}$ (defined using modulation spaces) is inspired from the weight class $D^{h}_{p}$ (defined using Lebesgue spaces, see Theorem 1.1). This provides the natural way to introduce weighted structures in modulation spaces, and turned out very effective in our analysis (see Remark 1.2 (2)).
@@ 2. Main results
Remark 2.2.
The hypothesis of Theorem 2.1 deserves several comments.
  1. (1)
    For the sufficiency part, the proof relies on maximal function techniques, where the comparison between the semigroup and the Hardy–Littlewood maximal operator requires $f\geq 0$ in order to apply pointwise domination arguments. See Remark 2.7 and the proof of Lemmas 4.1 and 4.2.
  2. (2)
    For the necessity part, the argument relies on the bounds of the heat kernel (and of the Hermite heat kernel) to derive the required lower estimates for the characterization of the weight class. In this direction, the non-negativity of $f$ is not required.
  3. (3)
    Due to certain additional algebraic properties of some modulation spaces, specifically when $1\leq p<\infty$ and $1\leq q\leq 2$, the non-negativity assumption can be relaxed. See Remark 2.9 and Theorem 4.6. In general, our method does not allow us to remove the non-negativity hypothesis. Whether the equivalence in Theorem 2.1 remains valid for general (possibly sign-changing) functions in $M^{p,q}_{v}$ is an interesting open question. See Section 7.
@@
    It is known (see e.g. [7, example 2.1]) that, for $0<\alpha<n,$ we have
@@
We shall now briefly comment on the proof of Theorem 2.1. These problems were motivated by the seminal work of Torrea et al. in [25], where they addressed similar issues in the setting of weighted Lebesgue spaces on $\mathbb{R}^{n}$. We observe that pointwise convergence results and the boundedness of the maximal operator are closely related. However, to the best of authors’ knowledge, it is important to note that the boundedness of the maximal operator on modulation spaces remains an open problem (Section 2.1), requiring the development of novel techniques, which may be of independent interest. Specifically, to prove Theorem 2.1, we employ various properties of modulation spaces to reduce the problem to a stage where we can apply known results [1, 25, 28] on weighted Lebesgue spaces. We refer to Remark 5.1 for detail proof strategy.
@@ 2.1. Maximal operator
Theorem 2.5 remains unknown for arbitrary functions (not necessarily non-negative) within the modulation spaces. This discrepancy arises from the fact that the definition of the maximal operator involves $|f|$, while certain modulation spaces do not necessarily remain closed under the modulus operation, see e.g. [10, Corollary 1.3] and [9, Corollary 3.2].
@@
It is worth noting that, by exploiting certain algebraic properties inherent to modulation spaces, Theorem 2.5 remains valid for arbitrary functions (not necessarily non-negative) belonging to specific modulation spaces, when $f\in M^{p,q}$ with $1\leq p<\infty$ and $1\leq q\leq 2$. Consequently, the pointwise convergence results stated in Theorem 2.1 also hold for such modulation spaces (see Theorem 4.6).
@@ 4. Hardy-Littlewood maximal operator on modulation spaces
In the definition of modulation spaces, since the choice of $\phi\in\mathcal{S}\setminus\{0\}$ is arbitrary, we may select $\phi$ such that $\phi(y)\leq 0$ for a.e. $y\in\mathbb{R}^{n}$, and ensure that $f*|\phi|\in L^{p}$ by utilizing the following chain of embeddings (see [43])
@@ 5. Heat equation with Laplacian
    By Lemma 5.2, the weight class on $\mathbb{R}^{2n}$, as considered in Theorem 2.1, can be suitably related to the previously studied weight class on $\mathbb{R}^{n}$ (see (1.8)). This correspondence ensures that the weight class on $\mathbb{R}^{2n}$ satisfies the necessary structural conditions required to apply existing results from the theory of weighted Lebesgue spaces.
@@
Hence, it follows that $g_{t_{0}}(x)<\infty$ a.e. for $x$, since we are given that
$\left\|h_{\frac{t_{0}}{8}}\right\|_{M_{v^{-1}}^{p^{\prime},q^{\prime}}}<\infty.$
Now, we will choose a weight $u\in L_{\text{loc}}^{1}$ such that $g_{t_{0}}\in M_{u}^{p,q}$, which is equivalent to showing $V_{\phi}g_{t_{0}}\in L_{u}^{p,q}$ for some $\phi\in\mathcal{S}.$ By choosing $u\in L_{\text{loc}}^{1}$ as
@@
Conversely, let
$\left\|h_{t_{0}}*f\right\|_{M_{u}^{p,q}}\lesssim\left\|f\right\|_{M_{v}^{p,q}}$
for all $f\in M_{v}^{p,q}$. Thus, $|h_{t_{0}}*f(x)|<\infty$ almost everywhere for all $f\in M_{v}^{p,q}$. Fix $x_{0}$ such that $h_{t_{0}}*f(x_{0})<\infty$. Then we will show that $h_{\frac{t_{0}}{4}}*f(x)<\infty$ for all $x$. Let assume $x\neq x_{0}$. Note that
@@ 6. Heat equation with Hermite operator
Consider the heat equation (1.1) associated with the harmonic oscillator $H$, whose heat semigroup is given by (1.5). The spectral decomposition of the Hermite operator is expressed in (1.4), where the eigenfunctions are the normalized Hermite functions
$\Phi_{\alpha}(x)=\prod_{j=1}^{n}\tilde{h}_{\alpha_{j}}(x_{j})$
and each $\tilde{h}_{k}$ denotes the one–dimensional normalized Hermite function, given by
$\tilde{h}_{k}(x)=\left(\sqrt{\pi},2^{k}k!\right)^{-1/2}(-1)^{k}e^{\frac{x^{2}}{2}}\frac{d^{k}}{dx^{k}}\big(e^{-x^{2}}\big).$
These functions form an orthonormal basis for $L^{2}$ and satisfy the eigenvalue relation
$H\Phi_{\alpha}=(2|\alpha|+n)\Phi_{\alpha}$
where $|\alpha|=\alpha_{1}+\dots+\alpha_{n}$. A fundamental identity involving Hermite functions is Mehler’s formula (see [42, Lemma 1.1.1]), which for $|w|<1$ takes the form
@@
Conversely, let us assume $\lim_{t\rightarrow 0}e^{-tH}f(x)=f(x)$ almost everywhere for all non-negative $f\in M_{v}^{p,q}$. Then, by (6.4) there exists $t_{0}>0$ such that
$e^{-tH}f(x)\lesssim h_{t_{0}}*f(x)<\infty$
almost everywhere for all non-negative $f\in M_{v}^{p,q}$. Following the same approach as in the converse part of Proposition 5.4, we can similarly deduce that $h_{\frac{t_{0}}{4}}\in M_{v^{-1}}^{p^{\prime},q^{\prime}}$, which implies $v\in D^{h}_{p,q}$. ∎
7. Concluding Remarks
Our main Theorems 2.1 and 2.5 naturally opens the door for the further investigation. In view of ongoing research interest mentioned in the introduction, we raise some interesting open questions in the following points.
  1. (1)
    Can we relax the the non-negativity assumption on the initial data in Theorem 2.1? The present proofs rely on positivity arguments and maximal function techniques. However, it would be of independent interest to determine whether the equivalence remains valid for general (possibly sign-changing) functions in weighted modulation spaces. See Remark 2.2.
  2. (2)
    It would be worthwhile to extend the analysis to other semigroups such as Poisson-type semigroups or semigroups generated by more general differential operators. See for e.g. the discussion following Theorem 1.1. In such settings, additional analytical difficulties may arise due to weaker kernel regularity or the absence of explicit representations.
  3. (3)
    Up to now our knowledge on maximal operator on modulation spaces is quite limited (see Theorem 2.5). The systematic investigation of the boundedness properties of maximal operators in modulation spaces remains an important open direction.
  4. (4)
    The combination of semigroup kernel estimates with time-frequency techniques provides a flexible method for deriving pointwise convergence theorems in non-classical function spaces. This shed new light on how large (rough) initial data one can take, see Remark 2.3. We expect that the ideas developed here may serve as a foundation for further results on almost everywhere convergence and related fine properties of other evolution equations in the modulation space setting.
@@ Acknowledgement
The second author gratefully acknowledges the support provided by IISER Pune, Government of India. We sincerely thank the referees for their careful reading of our manuscript and for their valuable comments and suggestions, which have significantly improved the presentation of the paper.